Excerpts from Candle Corporation's most recent balance sheet (in thousands of dollars) appear below:

Year 2 Year 1
Current assets:
Cash $ 168 $ 100
Accounts receivable, net 198 198
Inventory 148 138
Prepaid expenses 98 98
Total current assets $ 612 $ 534
Current liabilities:
Accounts payable $ 208 $ 188
Accrued liabilities 38 38
Notes payable, short term 98 88
Total current liabilities $ 344 $ 314

Sales on account during the year totaled $1,240 thousand. Cost of goods sold was $840 thousand.

Required:

Compute the following for Year 2:

a. Working capital. (Enter your answer in thousands of dollars, i.e., 100,000 should be entered as 100.)

b. Current ratio. (Round your answer to 2 decimal places.)

c. Acid-test (quick) ratio. (Round your answer to 2 decimal places.)

d. Accounts receivable turnover. (Round your answer to 2 decimal places.)

e. Average collection period. (Use 365 days in a year. Round your intermediate calculations to 2 decimals places and your final answer to 1 decimal place.)

f. Inventory turnover. (Round your answer to 2 decimal places.)

g. Average sale period. (Use 365 days in a year. Round your intermediate calculations to 2 decimals places and your final answer to 1 decimal place.)
Answer #1

a) Working capital = Current assets-Current liabilities = 612-344 = 268

b) Current ratio = Current assets/Current liabilities = 612/344 = 1.78

c) Acid test ratio = 366/344 = 1.06

d) Account receivable turnover = Sales/Average receivable = 1240/198 = 6.26

e) Average Collection period = 365/6.26 = 58.3 days

f) Inventory turnover = Cost of goods sold/Average inventory = 840/143 = 5.87

g) Average sale period = 365/5.87 = 62.2 days